OpenStudy (anonymous):

Can I please get help with this two part question. Part A: If (7^2)^x=1, what is the value of x? Part B: If (7^0)^x=1, what are the possible values for x?

OpenStudy (solomonzelman):

If you have \(\large\color{blue}{ a^x=1 }\) and \(\large\color{blue}{ a≠1 }\) then \(\large\color{blue}{ x=0 }\).

OpenStudy (anonymous):

So then part a would be 0?

OpenStudy (solomonzelman):

For the second one, when you have 1^x=1, x can be either ±1 or 0

OpenStudy (anonymous):

i cant see what u wrote it comes out as ??????

OpenStudy (solomonzelman):

yes part a is zero.

OpenStudy (solomonzelman):

refresh (?)

OpenStudy (solomonzelman):

you there?

OpenStudy (anonymous):

i keep on loosing connection

OpenStudy (solomonzelman):

happens

OpenStudy (anonymous):

solomonZelman do u mind helping me with a few more?

OpenStudy (solomonzelman):

sure, but do you know everything till now?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

alright

OpenStudy (anonymous):

(2^8 * 5^-5 * 19^0) * (5^-2)^4 * 2^28 ---- ( 2^3)

OpenStudy (anonymous):

should i draw it to make it easier to see?

OpenStudy (solomonzelman):

I see:)

OpenStudy (solomonzelman):

5^-2 will equal to what with a positive exponent ?

OpenStudy (anonymous):

1?

OpenStudy (solomonzelman):

\(\huge\color{blue}{ a^{-b}= \frac{1}{a^b} }\) For example. \(\huge\color{blue}{ 5^{-2}= \frac{1}{5^2}=\frac{1}{25} }\)

OpenStudy (solomonzelman):

So can you tell em what will 5^-5 be?

OpenStudy (solomonzelman):

don't simply just do the parts with positive exponents.

OpenStudy (anonymous):

so would it be 1/25?

OpenStudy (solomonzelman):

that is 5^-2. just leave it as 5^2 for now.

OpenStudy (solomonzelman):

I mean leave it as 1/5^2

OpenStudy (solomonzelman):

and 5^-5 is ?

OpenStudy (anonymous):

5*5*5*5*5?

OpenStudy (solomonzelman):

well, 5^-5 = 1/(5^5) right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So step 1: \(\huge\color{blue}{ \frac{(2^8 \times 5^{-5} \times 19^0)(5^{-2})^{4}\times 2^{28}}{2^3} }\) \(\huge\color{blue}{ \frac{(2^8 \times 5^{-5} \times 1)(5^{-2})^{4}\times 2^{28}}{2^3} }\) \(\huge\color{blue}{ \frac{(2^8 \times 5^{-5})(5^{-2})^{4}\times 2^{28}}{2^3} }\) good ?

OpenStudy (anonymous):

yes can you give me two minutes to write that down? plz

OpenStudy (solomonzelman):

sure

OpenStudy (solomonzelman):

Step 2: \(\huge\color{blue}{ \frac{(2^8 \times 5^{-5})(5^{-2})^{4}\times 2^{28}}{2^3} }\) \(\huge\color{blue}{ \frac{(2^8 \times 5^{-5})(5^{-8})\times 2^{28}}{2^3} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^8 \times 5^{-5} \times5^{-8}\times 2^{28}}{2^3} }\)

OpenStudy (anonymous):

in step 1 did u miss a number?

OpenStudy (solomonzelman):

I know that 19^0=1, right?

OpenStudy (solomonzelman):

and multiplying times 1, is just like not multiplying by anything.

OpenStudy (anonymous):

the first 1 in step 1?? theres no 2 to the 3rd power

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

can u give me two minutes plz to write it all down

OpenStudy (solomonzelman):

I can type it and you can copy it all.... or whatever works for you.

OpenStudy (anonymous):

i cant copy it, its not letitng me do that

OpenStudy (solomonzelman):

you can't copy these Latex symbols-:(

OpenStudy (solomonzelman):

you can write it on a paper, or take type to type it, it won't go anywhere.

OpenStudy (solomonzelman):

take time to type it *

OpenStudy (anonymous):

i typed it all

OpenStudy (solomonzelman):

so continuing from where I left off. The last thing I had is. \(\huge\color{blue}{ \frac{2^8 \times 5^{-5} \times5^{-8}\times 2^{28}}{2^3} }\)

OpenStudy (anonymous):

i got that down

OpenStudy (solomonzelman):

So, \(\huge\color{blue}{ \frac{2^8\times\frac{1}{5^5}\times\frac{1}{5^8}\times 2^{28}}{2^3} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^8\times 2^{28}}{2^3\times 5^5 \times 5^8} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^8\times 2^{28}}{2^3\times 5^{13}} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^3\times2^5\times 2^{28}}{2^3\times 5^{13}} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^5\times 2^{28}}{5^{13}} }\)

OpenStudy (solomonzelman):

\(\huge\color{blue}{ \frac{2^{33}}{5^{13}} }\)

OpenStudy (solomonzelman):

http://www.wolframalpha.com/input/?i=2%5E33%2F5%5E13

OpenStudy (solomonzelman):

you can write an answer in any form,. or just leave the last blue that I wrote. Depends on what your teacher exactly wants.

OpenStudy (anonymous):

ok im typing it give me two more minutes plz

OpenStudy (anonymous):

i got it thank u 1 more?

OpenStudy (anonymous):

The mass of the Eiffel Tower is about 9.16 ⋅ 10^6 kilograms. The mass of the Golden Gate Bridge is 8.05 ⋅ 10^8 kilograms. Approximately how many more kilograms is the mass of the Golden Gate Bridge than the mass of the Eiffel Tower? Show your work and write your answer in scientific notation.

OpenStudy (solomonzelman):

I am tired. Please close this question and open new one. You have an unlimited number of questions, but you have to ask 1 question per thread.

OpenStudy (solomonzelman):

tnx for your cooperation.

OpenStudy (anonymous):

ok so all i ogtta do is ask it on a new thread?

OpenStudy (anonymous):

thank u so so much for ur help